OpenStudy (anonymous):

15!/2!13! simplify answer

OpenStudy (shamil98):

Could you rephrase it without the !..

OpenStudy (shamil98):

It's hard to tell if that is 15/213 or 15/2 or 15i/213i

OpenStudy (tkhunny):

Really big hint: \(\dfrac{15!}{13!} = 15*14\)

OpenStudy (anonymous):

thats the way it is !

OpenStudy (shamil98):

o.o

hartnn (hartnn):

its called a factorial, and we use the property here that \(\large n! =n(n-1)!\)

OpenStudy (shamil98):

oh, well, i haven't learned that yet, my bad.

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

\[\frac{ 15! }{ 2!13! }=\frac{ 15\cdot14\cdot\cancel{13!} }{ 2!\cancel{13!} }\]

hartnn (hartnn):

so, \(\large 15! = 15\times 14! \\ \large 14! = 14 \times 13! \\ 2! =2*1\)

OpenStudy (anonymous):

can u simplify?

hartnn (hartnn):

after 13! gets cancelled, you are left with just 15*14/2 = ... ?

OpenStudy (anonymous):

?

hartnn (hartnn):

which step did u not understand ?

OpenStudy (anonymous):

so is15*14/2 the aswer?

hartnn (hartnn):

calculate it, notice that 14/2 = 7 so whats 15*7 =...?

OpenStudy (anonymous):

105

hartnn (hartnn):

yes, thats the answer.

OpenStudy (anonymous):

thanks a lot